OpenStudy (anonymous):

find the slope of the line through each pair of points (16,7) (-11,7)

OpenStudy (anonymous):

the slope = 0 lol

OpenStudy (anonymous):

@timo86m

OpenStudy (anonymous):

it will not let me message you

OpenStudy (anonymous):

use the slope formula m=(y2-y1)/(x2-x1)

OpenStudy (anonymous):

i did and it is = 0

OpenStudy (anonymous):

(16,7) (-11,7) (7-7)/(...) 0/any number = 0 So it is 0

OpenStudy (anonymous):

so do i get a medal to

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

timoo .. hoe about (16,-17)(-7,10)

OpenStudy (anonymous):

i dont get a medal =(

OpenStudy (anonymous):

(16,-17)(-7,10) (10-(-17))/(-7-16)= -1.17391

OpenStudy (anonymous):

27/-23=-(27/23)